OpenStudy (anonymous):

part A: solve A= 9/2(x+23) for x. Part B: determine the value of x when A= 108. part C: solve-np-90>30 for n. show your work

OpenStudy (anonymous):

\[A=\frac{9}{2}(x+23)\] like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1) multiply both sides by the reciprocal if \(\frac{9}{2}\) which is \(\frac{2}{9}\) when i say "multiply" i just mean write it \[\frac{2}{9}A=x+23\]

OpenStudy (anonymous):

last step is to subtract \(23\) from both sides

OpenStudy (anonymous):

how do I subtract from a fraction\[\frac{ 2 }{ 9}\]?